# RemixFusion: Residual-based Mixed Representation for Large-scale Online RGB-D Reconstruction

### v1

@@ 4.2. Quantitative and Qualitative Comparison
Figure 7 displays the reconstruction with estimated camera poses. Significant distortion is noticeable in the reconstructions produced by other methods, whereas our approach achieves accuracy and consistency. This underscores the precision of our pose estimation proposed in Section 3.3 and Section 3.4. Furthermore, the colorized trajectory in Figure 10 demonstrates the robustness and accuracy of our pose estimation even in very large-scale scenes, contrasting with failures or substantial drifts observed in alternative methods. We highly recommend readers refer to our supplementary video and materials for more comprehensive visualization and experimental results.
@@
Table 5 shows the quantitative results on room-level datasets, including Replica, ScanNet, and TUM RGB-D. Our method is the second-best on sequence scene000000 on ScanNet, which is challenging. Our method is comparable to the SOTA on average, and there is only a 0.6cm decrease in tracking accuracy compared to ESLAM. Our method performs better than ESLAM on Replica with 37.1 improvement. While ESLAM achieves the best performance on average, its system FPS is about 3, and does not meet the real-time requirements, which is important for the online SLAM system. Similar to ESLAM, LoopSplat showcases remarkable accuracy and robustness, which achieves the second-best tracking accuracy in this room-level benchmark thanks to the powerful loop closure modules. However, the modules are computationally expensive, and the efficiency is significantly limited (¡1 FPS). Our method achieves good tracking accuracy in large-scale scenes with relatively few iterations. Room-level scenarios require fewer optimizations, where we obtain comparable performance and great running efficiency (3 times faster than ESLAM and 21 times faster than LoopSplat) at the same time with the residual-based mixed representation. Therefore, our method is still superior to the others in terms of the online SLAM setting. Note that DROID-SLAM achieves higher FPS by prioritizing camera tracking over dense reconstruction. While the 3DGS-based methods like RTG-SLAM and GS-ICP SLAM are superior to other methods on Replica, they significantly fall short in real-world datasets like ScanNet and TUM RGB-D. This demonstrates the inherent instability of their camera tracking when exposed to real-world input noise. SplaTAM and MonoGS both use the rendering loss as the objective function, and demonstrate robustness in real-world room-level scenarios. Detailed comparisons of room-scale sequences on FastCaMo-Synth (noise-free), are provided in the supplementary materials.
@@
4.2.3. 3D Reconstruction.
The traditional explicit methods, including ElasticFusion and BundleFusion, can not succeed in finishing the tracking and reconstruction for almost all scenes. Therefore, we only report the results of the implicit methods, BAD-SLAM, and the explicit 3DGS-based method. Note that the reconstruction mesh for 3DGS-based methods (SplaTAM, MonoGS, RTG-SLAM, and GS-ICP SLAM) is obtained using TSDF Fusion with rendered RGB-D images following 2DGS (Huang et al., 2024b). Implicit methods often produce noisy meshes in empty spaces, posing challenges for fair comparisons. Following (Zhu et al., 2022; Wang et al., 2023), we use the estimated camera poses to cull the mesh with ground-truth depths for evaluation. Ground-truth meshes obtained from LiDAR scans use the same strategies with ground-truth poses. The evaluation is performed three times, and the average results are reported. Due to the scaling differences between the results of BAD-SLAM and ground-truth meshes, we align the output point clouds with ground-truth meshes before evaluation.
Figure 7 presents the reconstruction results, with an overview in the first row and a detailed zoom-in in the second row. Our methods can preserve the finest geometric details in these challenging large-scale scenes. In contrast, alternative methods struggle with unstable pose estimation and surface distortions, highlighting the effectiveness of our residual-based mixed representations. For example, in the second row of Figure 7, the comparisons of the white lines on the floor and black chairs illustrate that our reconstruction offers superior details in both texture and geometry.
Figure 8. Fully described in the text.
Figure 10 presents additional results about the large-scale scenes, which are challenging for both tracking and reconstruction. The colorized trajectory means the scanning timeline: red denotes the start and blue denotes the end. The area of dining on BS3D exceeds and there are multiple staircases and floors. Our method successfully finished the pose estimation and detailed reconstruction, whereas other approaches failed either in tracking or returning to the starting position, resulting in reconstruction distortions. For instance, MIPS-Fusion exhibits upward leaning and significant distortion in reconstruction upon returning to the second floor, despite employing loop closure for sub-maps. Note that the input depths for the floor of this scene are missing in the beginning, leading to some empty holes in the floor. The second column in Figure 10 shows the results of building2 on FastCaMo-Large, where there are three floors. Other methods notably fail on the third floor, where the right corridor should appear flat and straight. The third column in Figure 10 is a challenging self-captured sequence with faster camera motion, posing challenges for both tracking and mapping. RemixFusion demonstrates adaptability with robust pose estimation and efficient reconstruction, outperforming all the other methods. There are severe distortions for MIPS-Fusion at the end of the trajectory (blue), and the reconstruction is of great noise due to the heavy and poorly aligned sub-maps. Our reconstruction is clean and more accurate, with FPS 4 times faster than MIPS-Fusion. The other methods can succeed in reconstructing the first half of the sequence but fail in the middle, demonstrating the importance of robustness in large scenes.
@@
Photo-SLAM | 24.64 | 44.42 | 21.5 | 9.1G |
LoopSplat | 149.08 | 19.21 | 0.3 | 22.0G |
@@
GS-ICP SLAM is good at 2D photometric rendering, but its geometric rendering (D-L1) is notably inferior compared to other alternatives. This can be attributed to the reason that 3DGS-based representations achieve high-fidelity rendering through the fitting of 2D training views, indicating less attention to the 3D geometry in the online setting. This can also be proved by the metrics of 3D reconstruction accuracy and completeness presented in Table 4. In terms of real-time reconstruction, 3DGS-based methods are easy to achieve high-fidelity RGB rendering. However, they struggle to get accurate geometric reconstruction (see the D-L1 metric in Table 6). For example, RTG-SLAM performs well on the scene study of BS3D with 6.7cm in ATE RMSE (shown in Table 1). However, the 3D reconstruction quality (shown in Table 4) of it is not consistent with the performance of tracking. For comparison, Co-SLAM achieves 6.5cm on the scene study of BS3D in ATE RMSE, which is similar to RTG-SLAM; the 3D reconstruction of Co-SLAM is much better than RTG-SLAM regarding the accuracy and completeness. Note that RTG-SLAM focuses on the 3DGS optimization using the tracking module from other methods; its FPS and reconstruction are still inferior to Co-SLAM. Improving geometric reconstruction quality for 3DGS-based methods often requires more optimization iterations, which is generally unacceptable for real-time SLAM.
Furthermore, the comparison of novel view synthesis in Table 7 demonstrates that our residual-based representation significantly outperforms 3DGS in photometric rendering (PSNR) by 25 and surpasses it in geometric rendering (D-L1) by over 7 times. This indicates that the 3DGS-based methods are less geometry-aware in the online setting. Note that the sequence cafeteria-nvs of BS3D in Table 7 uses the manually partitioned train (50) and test images (50), with ground-truth poses excluding the effects of tracking. Overall, our residual-based representations surpass both implicit methods and explicit methods, including 3DGS, taking efficiency and performance into consideration in large scenes.
@@
Table 3 shows the rendering comparison on uHumans2 dataset. Our method outperforms the SOTA in terms of tracking by 12 and D-L1 by 12.3. Our method is the best in terms of PSNR on office, surpassing the SOTA by 0.33dB. While ESLAM is the second-best method for D-L1, its system FPS is over three times slower than our method. For 3DGS-based methods, MonoGS demonstrates the best RGB rendering on apartment (surpassing the second-best one by 4.49 for PSNR), but its depth rendering is far worse than other methods. This also indicates that 3DGS-based methods are less geometry-aware in the online setting, as mentioned above. Overall, our method is the best real-time method in terms of tracking and rendering on average.
@@ 4.2.5. Runtime and Memory Analysis
Table 8 illustrates the system FPS and the performance of both tracking and mapping. To ensure a comprehensive comparison, we focus on the scene lounge from BS3D, which is successfully handled by most methods. Note that the FPS of explicit methods like ElasticFusion and BundleFusion is high, but both of them fail in this scene. Therefore, they are not compared in this table. Our method achieves the best trade-off between GPU memory usage and accuracy. Note that our method maintains relatively consistent GPU memory usage regardless of scene scale, whereas explicit methods increase GPU memory usage with additional input frames in larger scenes. Thanks to the residual-based mapping, our method demonstrates superior system FPS compared to implicit methods, and the lightweight version of our method can run at 25 FPS with GB GPU memory usage, which is as fast as the explicit alternatives and requires less computational overhead.
Conversely, approaches like Co-SLAM and MIPS-Fusion employ hash schemes or multiple sub-maps to reduce memory consumption but struggle with efficiency and robustness in memorizing large-scale scenes. Methods like ESLAM, SplaTAM, MonoGS are not memory-friendly, requiring half-resolution images as input. Additionally, their system FPS is significantly lower than that of other methods. Similarly, LoopSplat is not memory-efficient in large scenes and exhibits running inefficiency (¡1 FPS), primarily due to the computationally expensive modules of loop closure and post-refinement. Photo-SLAM, leveraging ORB-SLAM3 for tracking and a map with hyper primitives for mapping, is lightweight and efficient. However, Photo-SLAM indicates worse tracking and reconstruction accuracy on large scenes. The traditional explicit method, like BAD-SLAM, is real-time but lacks robustness. GS-ICP SLAM, as the representative of 3DGS-based methods, is fast but compromises memory efficiency. Although RTG-SLAM leverages efficient techniques to reduce the redundant 3DGS and requires less GPU memory in the process, it requires large GPU memory for global optimization at the end of tracking. Our method stands out as the most efficient system, coupled with the most accurate pose estimation as well as reconstruction, considering both FPS and GPU memory usage.
@@ 4.3.1. Residual-based Mapping
One key innovation lies in the residual-based mapping, crucial for enabling real-time dense reconstruction of large-scale scenes in our system. To validate this design, we first compare our mixed method against using solely the implicit module without residuals. Parameters are the same except for the residual-based representations. Moreover, we demonstrate how the reconstruction quality varies according to the desired mapping FPS.
@@ 4.3.2. Residual-based bundle adjustment
Another key insight of our method is the residual-based bundle adjustment (RBA for short). Here we evaluate the effectiveness of our residual-based bundle adjustment for pose refinement. We compare our method to the traditional bundle adjustment that optimizes the individual pose variables (TBA for short). Additionally, we validate the effectiveness of our proposed gradient amplification (GA for short).
Compared to the commonly used bundle adjustment, which individually optimizes the pose of each frame, our proposed method employs a single MLP for 6D residual pose prediction, offering enhanced global consistency. Table 9 illustrates that the average ATE RMSE without bundle adjustment (baseline) is 6.39cm. Applying the RBA instead of TBA for pose refinement reduces the error marginally from 5.74cm to 5.68cm. Introducing GA to help bundle adjustment results in improved pose estimation for both TBA and RBA. Moreover, using RBA with GA reduces the error from 5.68cm to 4.61cm, resulting in an improvement of over 1cm. This proves the effectiveness of the proposed gradient amplification, which is simple yet necessary for the residual-based BA in large-scale scenes. Figure 14 demonstrates the comparison of different methods on foobar of BS3D, which validates the effectiveness and necessity of the proposed residual-based bundle adjustment and gradient amplification. In summary, the proposed residual-based BA can focus on detailed refinement of the initial camera poses and significantly improve the accuracy of pose estimation in large-scale scenes.
@@ 4.3.3. Adaptive Gradient Amplification
A key design in the residual-based BA is the proposed adaptive gradient amplification used in BA. As illustrated in Eq. 17, this technique amplifies the optimization gradients derived from the reconstructed surface, allowing the BA process to circumvent local minima and thus achieve more globally consistent solutions in real-world large-scale scenarios. The visualization of the optimization is shown in Figure 12. This design aims to mitigate the risk of falling into local optima, a common challenge in large-scale scenes. Visualization of the optimization of BA on waiting and corridor on BS3D is shown in Figure 12, where the 6D poses of all frames are visualized in 2D and colorized by the cost functions. Without the proposed gradient amplification, the optimization tends to get trapped in the local minimum (illustrated by the blue dashed line). However, the proposed gradient amplification enables BA to escape from the local minimum and obtain a better solution (orange dashed line).
Furthermore, we compare our method (GA) with the manually designed amplification (MA) and the randomized amplification (RA). Figure 13 presents bundle adjustment (BA) results under various truncation thresholds compared to the baseline (without GA). The positive improvement indicates the effectiveness of gradient amplification. Results of eight scenes of BS3D are reported as the range of the bar chart. GA achieves the best results on average with , showing positive improvements across all scenes. MA-5 and RA-5 indicate that the camera poses are guided towards the directions with 5cm, where the cameras are facing or randomized directions. Our proposed GA yields the most accurate pose estimation, while MA also improves the performance, which is less than GA. RA fails to improve the performance of bundle adjustment for all sequences. In summary, the improved results prove the effectiveness of the proposed gradient amplification.
@@
There is a primary limitation of our method. Robust pose estimation becomes challenging when a significant portion of the observed depth images is missing because our approach heavily depends on depth information. Consequently, our method can not be applied in scenarios where only RGB data is available. Failure cases of our method are shown in Figure 15. The majority of observed depth information in the middle of the sequence (highlighted by red rectangles) is missing, making it hard and unstable for both pose estimation and reconstruction.

### v3

@@ 4.2. Quantitative and Qualitative Comparison
Figure 7 displays the reconstruction with estimated camera poses. Significant distortion is noticeable in the reconstructions of other methods, whereas our approach delivers superior accuracy and consistency. This underscores the precision of our pose estimation proposed in Section 3.3 and Section 3.4. Furthermore, the colorized trajectory in Figure 10 demonstrates the robustness and accuracy of our pose estimation even in very large-scale scenes, contrasting with failures or substantial drifts observed in other methods. We highly recommend readers refer to our supplementary video and materials for more comprehensive visualization.
@@
Table 5 shows the quantitative results on room-level datasets, including Replica, ScanNet, and TUM RGB-D. Our method is the second-best on scene000000 of ScanNet, which is challenging. Our method is comparable to the SOTA on average, and there is only a 0.6cm decrease in tracking accuracy compared to ESLAM and LoopSplat. Our method performs better than ESLAM on Replica with 37.1 improvement. While ESLAM achieves the second-best performance on average, its system FPS is about 3, and does not meet the real-time requirements, which is important for the online SLAM system. Similar to ESLAM, LoopSplat showcases remarkable accuracy and robustness, which achieves the best tracking accuracy in this room-level benchmark thanks to the powerful loop closure modules. However, the modules are computationally expensive, and the efficiency is significantly limited (¡1 FPS). Room-level scenarios require fewer optimizations, where we obtain comparable performance and great running efficiency (3 times faster than ESLAM and 21 times faster than LoopSplat) at the same time with the residual-based mixed representation. Overall, our method is still superior to the others in terms of the online SLAM setting. Note that DROID-SLAM achieves higher FPS by prioritizing camera tracking over dense reconstruction. While the 3DGS-based methods like RTG-SLAM and GS-ICP SLAM are superior to other methods on Replica, they significantly fall short in real-world datasets like ScanNet and TUM RGB-D. This demonstrates the inherent instability of their camera tracking when exposed to real-world input noise. SplaTAM and MonoGS both use the rendering loss as the objective function, and demonstrate robustness in real-world room-level scenarios. Detailed comparisons of room-scale sequences on FastCaMo-Synth (noise-free), are provided in the supplementary materials.
@@
4.2.3. 3D Reconstruction.
The traditional explicit methods, including ElasticFusion and BundleFusion, can not succeed in finishing the tracking and reconstruction for almost all scenes. Therefore, we only report the results of the implicit methods, BAD-SLAM, and the explicit 3DGS-based method. Note that the reconstruction mesh for 3DGS-based methods is obtained using TSDF Fusion with rendered RGB-D images following 2DGS (Huang et al., 2024b). Implicit methods often produce noisy meshes in empty spaces, posing challenges for fair comparisons. Following (Zhu et al., 2022; Wang et al., 2023), we use the estimated camera poses to cull the mesh with ground-truth depths for evaluation. Ground-truth meshes obtained from LiDAR scans use the same strategies with ground-truth poses. The evaluation is performed three times, and the average results are reported. Due to the scaling differences between the results of BAD-SLAM and ground-truth meshes, we align the output point clouds with ground-truth meshes before evaluation.
Figure 7 presents the reconstruction results, with an overview in the first row and a detailed zoom-in in the second row. Our methods can preserve the finest geometric details in these challenging large-scale scenes. In contrast, alternative methods struggle with unstable pose estimation and surface distortions, highlighting the effectiveness of our residual-based mixed representations. For example, in the second row of Figure 7, the comparisons of the white lines on the floor and black sofas illustrate that our reconstruction offers superior details in both texture and geometry.
Figure 8. Fully described in the text.
@@
Figure 10 presents additional results about the large-scale scenes, which are challenging for both tracking and reconstruction. The colorized trajectory means the scanning timeline: red denotes the start and blue denotes the end. The area of dining on BS3D exceeds and there are multiple staircases and floors. Our method successfully finished the pose estimation and detailed reconstruction, whereas other approaches failed either in tracking or returning to the starting position, resulting in reconstruction distortions. For instance, MIPS-Fusion exhibits upward leaning and significant distortion in reconstruction upon returning to the second floor, despite employing loop closure for sub-maps. Note that the input depths for the floor of this scene are missing in the beginning, leading to some empty holes in the reconstruction. The second column in Figure 10 shows the results of building2 on FastCaMo-Large, where there are three floors. Other methods notably fail on the third floor, where the right corridor should appear flat and straight. The third column in Figure 10 is a challenging self-captured sequence with faster camera motion, posing challenges for both tracking and mapping. RemixFusion demonstrates adaptability with robust pose estimation and efficient reconstruction, outperforming all the other methods. There are severe distortions for MIPS-Fusion at the end of the trajectory (blue), and the reconstruction is of great noise due to the heavy and poorly aligned sub-maps. Our reconstruction is clean and more accurate, with FPS 4 times faster than MIPS-Fusion. The other methods can succeed in reconstructing the first half of the sequence but fail in the middle, demonstrating the importance of robustness in large scenes.
@@
| Photo-SLAM | 24.6 | 44.42 | 21.5 | 9.1G |
| LoopSplat | 149.1 | 19.21 | 0.3 | 22.0G |
@@
GS-ICP SLAM is good at 2D photometric rendering, but its geometric rendering (D-L1) is notably inferior compared to other alternatives. This can be attributed to the reason that 3DGS-based representations achieve high-fidelity rendering through the fitting of 2D training views, indicating less attention to the 3D geometry in the online setting. This can also be proved by the metrics of 3D reconstruction accuracy and completeness presented in Table 4. In terms of real-time reconstruction, 3DGS-based methods are easy to achieve high-fidelity RGB rendering. However, they struggle to get accurate geometric reconstruction (see the D-L1 metric in Table 6). For example, RTG-SLAM performs well on the scene study of BS3D with 6.7cm in ATE RMSE (shown in Table 1). However, its 3D reconstruction quality (shown in Table 4) is not consistent with the performance of tracking. For comparison, Co-SLAM achieves 6.5cm on the scene study of BS3D in ATE RMSE, which is similar to RTG-SLAM; the 3D reconstruction of Co-SLAM is much better than RTG-SLAM regarding the accuracy and completeness. Note that RTG-SLAM focuses on the 3DGS optimization using the tracking module from other methods. Its FPS and reconstruction are still inferior to Co-SLAM. Improving geometric reconstruction quality for 3DGS-based methods often requires more optimization iterations, which is generally unacceptable for real-time SLAM.
Furthermore, the comparison of novel view synthesis in Table 7 demonstrates that our residual-based representation significantly outperforms 3DGS in photometric rendering (PSNR) by 25 and surpasses it in geometric rendering (D-L1) by over 7 times. This indicates that the 3DGS-based methods are less geometry-aware in the online setting. The difference between training views and novel views indicates the robustness of our mixed representation. Note that the sequence cafeteria-nvs of BS3D in Table 7 uses the manually partitioned train (50) and test images (50), with ground-truth poses excluding the effects of tracking. Overall, our residual-based representations surpass both implicit methods and explicit methods, including 3DGS, taking efficiency and performance into consideration in large scenes.
@@
Table 3 shows the rendering comparison on uHumans2 dataset. Our method outperforms the SOTA in terms of tracking by 12 and D-L1 by 12.3. Our method is the best in terms of PSNR on office, surpassing the SOTA by 0.34dB. While ESLAM is the second-best method for D-L1, its system FPS is over three times slower than our method. For 3DGS-based methods, MonoGS demonstrates the best RGB rendering on apartment (surpassing the second-best one by 4.49dB for PSNR), but its depth rendering is far worse than other methods. This also indicates that 3DGS-based methods are less geometry-aware in the online setting, as mentioned above. Overall, our method is the best real-time method in terms of tracking and rendering on average.
@@ 4.2.5. Runtime and Memory Analysis
Table 8 illustrates the system FPS and the performance of both tracking and mapping. To ensure a comprehensive comparison, we focus on the scene lounge from BS3D, which is successfully handled by most methods. Note that the FPS of explicit methods like ElasticFusion and BundleFusion is high, but both of them fail in this scene. Therefore, they are not included in this table. Our method achieves the best trade-off between GPU memory usage and accuracy. Our method exhibits relatively consistent GPU memory usage across different scene scales. In contrast, the memory consumption of explicit methods grows substantially as they process more frames in larger scenes. Our mixed representation demonstrates superior system FPS compared to implicit methods, and the lightweight version of our method can run at 25 FPS with GB GPU memory usage, which is as fast as the explicit alternatives and requires less computational overhead.
Conversely, approaches like Co-SLAM and MIPS-Fusion employ hash schemes or multiple sub-maps to reduce memory consumption but struggle with efficiency and robustness in memorizing large-scale scenes. Methods like ESLAM, SplaTAM, and MonoGS are not memory-friendly, requiring half-resolution images as input. Additionally, their system FPS is significantly lower than that of other methods. Similarly, LoopSplat is not memory-efficient in large scenes and exhibits running inefficiency (¡1 FPS), primarily due to the computationally expensive modules of loop closure and post-refinement. Photo-SLAM, leveraging ORB-SLAM3 for tracking and a map with hyper primitives for mapping, is lightweight and efficient. However, Photo-SLAM indicates worse tracking and reconstruction accuracy on large scenes. The traditional explicit method, like BAD-SLAM, is real-time but lacks robustness. GS-ICP SLAM, as the representative of 3DGS-based methods, is fast but compromises memory efficiency. Although RTG-SLAM leverages efficient techniques to reduce the redundant 3DGS and requires less GPU memory in the process, it requires large GPU memory for global optimization at the end of tracking. Our method stands out as the most efficient system, coupled with the most accurate pose estimation as well as reconstruction, considering both FPS and GPU memory usage.
@@ 4.3.1. Residual-based Mapping
One key innovation is the residual-based mapping, crucial for enabling real-time dense reconstruction of large-scale scenes in our system. To validate this design, we first compare our mixed method against using solely the implicit module without residuals. Parameters are the same except for the residual-based representations. Moreover, we demonstrate how the reconstruction quality varies according to the desired mapping FPS.
@@ 4.3.2. Residual-based bundle adjustment
Another key insight of our method is the residual-based bundle adjustment (RBA for short). Here we evaluate the effectiveness of our residual-based bundle adjustment for pose refinement. We compare our method to the traditional bundle adjustment (TBA for short) that optimizes the individual pose variables. Additionally, we validate the effectiveness of our proposed gradient amplification (GA for short).
@@
Compared to the commonly used bundle adjustment, which individually optimizes the pose of each frame, our proposed method employs a single MLP for 6D residual pose prediction, offering enhanced global consistency. Table 9 illustrates that the average ATE RMSE without bundle adjustment (baseline) is 6.39cm. Applying the RBA instead of TBA for pose refinement reduces the error marginally from 5.74cm to 5.68cm. Introducing GA to help bundle adjustment results in improved pose estimation for both TBA and RBA. Moreover, using RBA with GA reduces the error from 5.68cm to 4.61cm, resulting in an improvement of over 1cm. This proves the effectiveness of the proposed gradient amplification, which is simple yet effective for the residual-based BA in large-scale scenes. Figure 14 demonstrates the comparison of different methods on foobar of BS3D, which validates the effectiveness and necessity of the proposed residual-based bundle adjustment and gradient amplification. In summary, the proposed residual-based BA can focus on detailed refinement of the initial camera poses and significantly improve the accuracy of pose estimation in large-scale scenes.
@@ 4.3.3. Adaptive Gradient Amplification
A key design in the residual-based BA is the proposed adaptive gradient amplification used in BA. As illustrated in Eq. 17, this technique amplifies the optimization gradients derived from the reconstructed surface, allowing the BA to circumvent local minima and thus achieve more globally consistent solutions in real-world large-scale scenarios. This design aims to mitigate the risk of falling into local optima, a common challenge in large-scale scenes. Visualization of the optimization of BA on waiting and corridor of BS3D is shown in Figure 12, where the 6D poses of all keyframes are visualized in 2D and colorized by the cost functions. Without the proposed gradient amplification, the optimization tends to get trapped in the local minimum (illustrated by the blue dashed line). However, the proposed gradient amplification enables BA to escape from the local minimum and obtain a better solution (the orange dashed line).
Furthermore, we compare our method (GA) with the manually designed amplification (MA) and the randomized amplification (RA). Figure 13 presents the improvement of BA under various truncation thresholds compared to the baseline (without GA). The positive improvement indicates the effectiveness of gradient amplification. Results of eight scenes of BS3D are reported as the range of the bar chart, and the red line indicates the median. GA achieves the best results on average with , showing positive improvements across all scenes. MA-5 and RA-5 indicate that the camera poses are guided towards the directions with 5cm, where the cameras are facing or randomized directions. Our proposed GA yields the most accurate pose estimation, while MA also improves the performance, which is less than GA. RA fails to improve the performance of bundle adjustment for all sequences. In summary, the improved results prove the effectiveness of the proposed gradient amplification.
@@
There is a primary limitation of our method. Robust pose estimation becomes challenging when the depth images contain extensive missing regions, as our approach critically relies on depth information. Consequently, our method can not be applied in scenarios where only RGB images are available. Another limitation is that our method is not robust in dynamic environments. Failure cases of our method are shown in Figure 15. The majority of observed depth information in the middle of the sequence (highlighted by red rectangles) is missing, making it hard and unstable for both pose estimation and reconstruction.